Find 3 consecutive numbers such that the sum of the first and three times the second is 44 less than 5 times the third
Answer by boilpoil(127)

Let x be the first number
x+3(x+1)=5(x+2)-44
x+3x+3=5x+10-44
4x+3=5x-34
-x=-37
x=37
The second number
37+1
=38
The third number
37+2
=39